A candy store sells cans of cashews for $4.50 per pound and peanuts for $2.50 per pound. If you wanted to buy exactly 8 pounds of nuts for $23.00, how many pounds of each kind should you buy?
Answer by stanbon(75887)

A candy store sells cans of cashews for $4.50 per pound and peanuts for $2.50 per pound. If you wanted to buy exactly 8 pounds of nuts for $23.00, how many pounds of each kind should you buy?
--------------------------
Note: $23 for 8 lbs means 2.875 per pound
-----------------------------------
Value Equation: 4.5c + 2.5p = 2.875*8
Quantity Equation: c + p = 8 lbs
----------------------
c = 8-p
----
Substitute for "c" and solve for "p":
4.5(8-p) + 2.5p = 2.875*8
---
4.5*8 - 4.5p + 2.5p = 2.875*8
-----
-2.5p = -1.625*8
p = 5.2 lbs (amt. of peanuts needed)
----
c = 8-5.2 = 2.7 lbs (amt. of cashews needed)
===============================================
Cheers,
Stan H.
==========